craig is 28 years younger than his father the sum of both ages is 64 how old is craigs father

Answer by Maths68(1474)

Let
Present age of Craig = c
Present age of Father = f
Craig is 28 years younger means if we add 28 to the present age of Craig, both father and son will be of equal age
f=c+28 ........(1)
Sum of their ages
c+f=64.........(2)
Put the value of f from (1) to (2)
c+(c+28)=64
c+c+28=64
2c+28=64
2c=36
Divide by 2 both sides of above equation
2c/2=36/2
c=18
Put the value of c in (1)
f=c+28 ........(1)
f=18+28
f=46
Present age of Craig = c = 18 years old
Present age of Father = f = 46 years old